# Low-Rank Interconnected Adaptation across Layers

### v2

@@ Appendix E Where to Apply Lily in Mamba?
where $h_{t}$ is the hidden state at time $t$ and $x_{t}$ is the corresponding input token. Delta projection is a module in SSM that’s learnable and tasked with transforming the parameter $\Delta$. Since adapting the delta projection alone can indirectly adapt the entire SSM module (i.e., $\bar{A}$ and $\bar{B}$ are all determined by $\Delta$), it is the most critical component of the SSM module.
We investigate the performance of two adaptation strategies: adapting only the input linear projection layer (denoted as "in") and adapting both the input linear projection layer and SSM (denoted as "$\Delta$ + in" since we only adapt delta projection in SSM module). Our results, as shown in Table 1, indicate that applying Lily solely to the input projection yields better performance than applying it to both the input and delta projection modules. This suggests that when adapting Mamba-based models under the paradigm of low-rank adaptation, it is optimal to adapt only the input projection module outside the SSM module. These findings highlight the need for further research into the impact of fine-tuned modules in Mamba on overall performance. Additionally, developing PEFT methods specifically tailored to Mamba-based models, whether for vision or language foundation models, is also a promising direction for future work.
Appendix F Performance with Different Learning Rates
@@
Appendix G Does Selectivity Help?
@@
As shown in Table 17, removing selectivity from Lily results in generally poorer performance compared to vanilla Lily. This is likely because the lack of selectivity causes Lily to simply aggregate all the HP expert, leading to inferior performance. This validates the design choice of using routers in Lily to selectively allocate weights to HP experts, rather than simply summing them.
@@
Appendix H How to Allocate Parameters?
Since Lily alters the traditional LoRA’s layer-bound setup, increasing the parameters of Lily can be achieved through two approaches: 1) increasing ne, i.e., increasing the number of LP and HP experts, and 2) increasing the rank, i.e., increasing the parameter size of each individual LP or HP expert. In this section, we investigate which factor has the greatest impact on performance. We conduct experiments on the commonsense reasoning task. Specifically, we maintain the same parameter count and learning rate, and achieve the same parameter count by setting different ranks and adjusting the corresponding ne (e.g., r=16, ne=4 versus r=8, ne=8). The results are shown in Fig. 9, from which we observe that more LP and HP experts with smaller rank (i.e., bigger ne and smaller rank) generally performs worse. We argue that this is because, although increasing the attention granularity allows for finer details, the resulting performance gain is not as significant as the gain obtained by increasing the rank, i.e., increasing the model’s capacity to learn more information. This gives us an insight that, in Lily, increasing ne to increase the parameters is less effective than directly increasing the rank in terms of potential performance gain.
@@ Appendix H How to Allocate Parameters?
Appendix I More on Subject-driven Generation
@@
Appendix J More on Attention Maps of Lily and LoRA

### v3

@@ Appendix F Where to Apply Lily in Mamba?
where $h_{t}$ is the hidden state at time $t$ and $x_{t}$ is the corresponding input token. Delta projection is a module in SSM that’s learnable and tasked with transforming the parameter $\Delta$. Since adapting the delta projection alone can indirectly adapt the entire SSM module (i.e., $\bar{A}$ and $\bar{B}$ are determined by $\Delta$), it is the most critical component of the SSM module.
We investigate the performance of two adaptation strategies: adapting only the input linear projection layer (denoted as "in") and adapting both the input linear projection layer and the SSM (denoted as "$\Delta$ + in" since we only adapt the delta projection in the SSM module). Our results, as shown in Table 1, indicate that applying Lily solely to the input projection yields better performance than applying it to both the input and delta projection modules. This suggests that when adapting Mamba-based models under the paradigm of low-rank adaptation, it is optimal to adapt only the input projection module outside the SSM module. These findings highlight the need for further research into the impact of fine-tuned modules in Mamba on overall performance. Additionally, developing PEFT methods specifically tailored to Mamba-based models, whether for vision or language foundation models, is also a promising direction for future work.
Appendix G Performance with Different Learning Rates
@@
Appendix H Does Selectivity Help?
@@
As shown in Table 17, removing selectivity from Lily results in generally poorer performance compared to vanilla Lily. This is likely because the lack of selectivity causes Lily to simply aggregate all the $B$ expert, leading to inferior performance. This validates the design choice of using routers in Lily to selectively allocate weights to $B$ experts, rather than simply summing them.
@@
Appendix I How to Allocate Parameters?
Since Lily alters the traditional LoRA’s layer-bound setup, increasing the parameters of Lily can be achieved through two approaches: 1) increasing ne, i.e., increasing the number of $A$ and $B$ experts, and 2) increasing the rank, i.e., increasing the parameter size of each individual $A$ or $B$ expert. In this section, we investigate which factor has the greatest impact on performance. We conduct experiments on the commonsense reasoning task. Specifically, we maintain the same parameter count and learning rate, and achieve the same parameter count by setting different ranks and adjusting the corresponding ne (e.g., r=16, ne=4 versus r=8, ne=8). The results are shown in Fig. 9, from which we observe that more $A$ and $B$ experts with smaller rank (i.e., bigger ne and smaller rank) generally performs worse. We argue that this is because, although increasing the attention granularity allows for finer details, the resulting performance gain is not as significant as the gain obtained by increasing the rank, i.e., increasing the model’s capacity to learn more information. This gives us an insight that, in Lily, increasing ne to increase the parameters is less effective than directly increasing the rank in terms of potential performance gain.
@@ Appendix I How to Allocate Parameters?
Appendix J More on Subject-driven Generation
@@
Appendix K From a Feature Merging Perspective
Apart from having higher-rank weight updates than LoRA, Lily also enables comprehensive information access across layers. Lily enables access to information or features from all other layers when adapting a target module at a specific layer thanks to the inter-connectivity of the adapters. We aim to understand how Lily achieves this comprehensive information access from the perspective of visual tasks as shown in Fig. 12. We can observe that, in Lily, the distinctness of the attention maps between layers is not as pronounced as in LoRA. This validates Lily’s ability to enable all-level information access, since adaptation at each layer takes into account features from other layers. Additionally, we specifically visualize the actual feature differences between different layers in Fig. 13. We observe that Lily has more points with low feature differences (blue color) than LoRA, indicating that the distinctness of features between layers in Lily is generally lower than in LoRA. This further demonstrates Lily’s ability to enable comprehensive information access. Although we enable all-level information access, what prevents the features from becoming completely identical is the selectivity introduced by Lily, which we specify in the following section.
Appendix L More on Attention Maps of Lily and LoRA